# Technical Report of HelixFold3 for Biomolecular Structure Prediction

### v2

@@ Abstract
To address these challenges, the PaddleHelix team is developing HelixFold3, aiming to replicate AlphaFold3’s capabilities. Leveraging insights from previous models and extensive datasets, HelixFold3 achieves accuracy comparable to AlphaFold3 in predicting the structures of the conventional ligands, nucleic acids, and proteins. The initial release of HelixFold3 is available as open source on GitHub for academic research, promising to advance biomolecular research and accelerate discoveries.
@@ 1 Introduction
The initial release of HelixFold3, which includes the inference code and the current version of model parameters, is now available as open source on PaddleHelix’s GitHub repository. You can access it at https://github.com/PaddlePaddle/PaddleHelix/blob/dev/apps/protein_folding/helixfold3 for academic research. This release is intended for non-commercial use and provides researchers with the tools needed to explore and leverage HelixFold3’s advanced capabilities in biomolecular structure prediction. We believe that the open-source availability of HelixFold3 will significantly contribute to the advancement of research in biomolecular interactions. We also provide online service at PaddleHelix website at https://paddlehelix.baidu.com/app/all/helixfold3/forecast. Researchers can now build upon this foundation, conduct further studies, and apply HelixFold3 to a broader range of biological questions, thereby advancing our understanding of complex biomolecular systems and accelerating the development of new applications in structural biology and related areas.
@@ 2 Results
We begin by evaluating the performance of HelixFold3 in multiple datasets. For ligands, we utilize the PoseBusters benchmark [9] to assess precision and physical plausibility. For nucleic acids, HelixFold3 is evaluated on CASP15 [10] RNA targets and recent RNA and DNA structures from the RCSB Protein Data Bank (RCSB PDB) [8]. We also assessed the model’s precision for protein structure predictions on CASP15 protein targets. Each sample is processed with 5 different random seeds, and diffusion inference is performed 5 times per seed with 200 sampling steps. The most accurate predictions are selected based on their confidence scores.
@@ 2.2 Nucleic Acids
Accurately predicting the structure of nucleic acid targets in a fully automated manner, without human input, remains a formidable challenge, primarily due to the limited availability of crystallized nucleic acid structures. To evaluate HelixFold3’s capabilities in this domain, we conducted a comparative analysis against several baseline methods, starting with RNA targets from CASP15 [10], using the evaluation framework established by AlphaFold3. Figure 2(a) illustrate a comparison between HelixFold3 and the baseline methods, showcasing the average RNA LDDT across 12 targets as well as the RNA LDDT for each individual target. While HelixFold3’s accuracy on CASP15 RNA samples does not yet match that of AIchemy_RNA2 [11], which benefits from human intervention, it achieves a level of accuracy on par with AlphaFold3 among models that operate in a fully automated manner.
@@ 2.3 Proteins
For protein-protein complex structure prediction, AlphaFold-Multimer represents a significant advancement over earlier models, though its success rate and accuracy still have room for improvement. AlphaFold3 further enhances these capabilities, delivering superior predictive performance. We analyzed 186 protein complexes released in the PDB from January 19, 2022, to November 30, 2022 to assess HelixFold3 and the competitive methods. As depicted in Figure 3, HelixFold3 has already outperformed AlphaFold-Multimer in predicting protein-protein interfaces, yet a gap remains between HelixFold3 and AlphaFold3. To address this, ongoing research will concentrate on targeted optimizations and iterative refinements of HelixFold3, with the aim of achieving greater accuracy and reliability in protein-protein complex predictions.
@@
(b) Success rate of protein-protein interface targets collected from PDB

### v3

@@ Abstract
To address these challenges, the PaddleHelix team is developing HelixFold3, aiming to replicate AlphaFold3’s capabilities. Leveraging insights from previous models and extensive datasets, HelixFold3 achieves accuracy comparable to AlphaFold3 in predicting the structures of the conventional ligands, nucleic acids, and proteins. The initial release of HelixFold3 is available as open source on GitHub for academic research, promising to advance biomolecular research and accelerate discoveries. The latest version will be continuously updated on the HelixFold3 web server, providing both interactive visualization and API access.
@@ 1 Introduction
The initial release of HelixFold3, which includes the inference code and the current version of model parameters, is now available as open source on PaddleHelix’s GitHub repository. You can access it at https://github.com/PaddlePaddle/PaddleHelix/blob/dev/apps/protein_folding/helixfold3 for academic research. This release is intended for non-commercial use and provides researchers with the tools needed to explore and leverage HelixFold3’s advanced capabilities in biomolecular structure prediction. We believe that the open-source availability of HelixFold3 will significantly contribute to the advancement of research in biomolecular interactions. We also provide online service at PaddleHelix website at https://paddlehelix.baidu.com/app/all/helixfold3/forecast, providing both interactive visualization and API access. HelixFold3 also provides a simple and user-friendly API service (). Users can easily schedule Baidu Intelligent Cloud resources with just a few simple steps, enabling rapid execution of tens of thousands of structure predictions without the need for expensive hardware or complex configurations. The API supports seamless integration with existing bioinformatics toolchains and is widely applicable across various fields of scientific research and commercial applications.
@@ 2 Results
Due to ongoing iterations of HelixFold3, we have labeled the version released on GitHub in August 2024 as HelixFold3, and the version released in December 2024 on the PaddleHelix web server as HelixFold3.1 in subsequent results.
We begin by evaluating the performance of HelixFold3 in multiple datasets. For ligands, we utilize the PoseBusters benchmark [9] to assess precision and physical plausibility. For nucleic acids, HelixFold3 is evaluated on CASP15 [10] RNA targets and recent RNA, DNA, and protein-RNA complex structures from the RCSB Protein Data Bank (RCSB PDB) [8]. We also evaluated the model’s precision in predicting protein structures using recently released protein-protein complexes from the PDB and the SAbDab database [11], which serves as an antigen-antibody evaluation set. Each sample is processed with 5 different random seeds, and diffusion inference is performed 5 times per seed with 200 sampling steps. The prediction with the highest confidence score was used for the evaluation.
@@ 2.2 Nucleic Acids
Accurately predicting the structure of nucleic acid targets in a fully automated manner, without human input, remains a formidable challenge, primarily due to the limited availability of crystallized nucleic acid structures. To evaluate HelixFold3’s capabilities in this domain, we conducted a comparative analysis against several baseline methods, starting with RNA targets from CASP15 [10], using the evaluation framework established by AlphaFold3. Figure 2(a) illustrates a comparison between HelixFold3 and the baseline methods, showcasing the average RNA LDDT across 12 targets as well as the RNA LDDT for each individual target. HelixFold3.1 outperforms other fully automated models and stands on par with AIchemy_RNA2 [12], which benefits from human intervention.
@@
We also evaluate the accuracy of HelixFold3 on the protein-nucleic acid complexes, using the iLDDT metric within a 30Å range for local structures. As illustrated in Figure 2(c), HelixFold3 performs comparable to AlphaFold3 and significantly outperforms RF2NA in predicting protein-RNA and protein-double-stranded DNA complexes. Furthermore, HelixFold3 slightly outperforms AlphaFold3 in protein-RNA complexes.Note that the results for AlphaFold3 are taken directly from the AlphaFold3 paper, making the comparison not entirely fair.
@@ 2.3 Proteins
For protein-protein complex structure prediction, AlphaFold-Multimer represents a significant advancement over earlier models, though its success rate and accuracy still have room for improvement. AlphaFold3 further enhances these capabilities, delivering superior predictive performance. We analyzed 186 protein complexes released in the PDB from January 19, 2022, to November 30, 2022, to assess HelixFold3 and the competitive methods. As depicted in Figure 3, HelixFold3 has already outperformed AlphaFold-Multimer in predicting protein-protein interfaces, yet a gap remains between HelixFold3 and AlphaFold3. To address this, ongoing research will concentrate on targeted optimizations and iterative refinements of HelixFold3, with the aim of achieving greater accuracy and reliability in protein-protein complex predictions.
[…]
@@
(b) Success rate of protein-protein interface targets collected from PDB.
(c) Prediction precise of antigen-antibody collected from SAbDab[11].
(d) Prediction precise of antigen-antibody with different numbers of specified epitope residues
@@
2.4 Covalent Modification
Figure 4: Results for covalent modification
Covalent modifications are essential for the regulation of protein function, stability, and interactions, yet predicting their structural impact remains a significant challenge for computational methods. Following the evaluation protocol outlined in the AlphaFold3 paper, MMseqs was used to cluster covalently modified structures sourced from RecentPDB. It should be noted that our cluster counts differ slightly from those reported in the AlphaFold3 paper, probably due to minor differences in implementation details or database versions. As a result, direct comparisons of performance may not be entirely fair. As shown in Figure 4, HelixFold3 demonstrates a higher precision than AlphaFold3 in predicting high quality single-residue glycosylation and RNA modification structures, measured by the percentage of successful predictions with RMSD <2Å. For other modifications, both models exhibit comparable performance.